OpenStudy (anonymous):

A projectile is thrown upward so that its distance above the ground after t seconds is h = -16t2 + 672t. After how many seconds does it reach its maximum height?

OpenStudy (jdoe0001):

so you have h = -16t2 + 672t which is really a parabola, with a negative leading coefficient, meaning is a parabola going downwards if we factor out the parabola form, to a vertex form of a parabola, we'd get the "vertex" of it, that is, its (h, k) or (x, y) values how many "x" seconds it took to get to "y" height

OpenStudy (anonymous):

that still doesn't make sense .-.

OpenStudy (jdoe0001):

so let's start by getting common factor off it h(t) = -16t2 + 672t => \(h(t) -16(t^2-42t)\) now from there, we'd need a number to complete the "perfect trinomial" that is a number at \(\large h(t) -16(t^2-42t+\square?)\)

OpenStudy (jdoe0001):

keep in mind that, to complete the square, all we're doing is using Mr zero 0 :) that is, if we add say 200, we must also substract 200 so if we borrow 200, we must give back 200 to get 0 back

OpenStudy (jdoe0001):

any number inside the parentheses, will be multiplied by -16, as you can see, thus the number will end up being a negative number, so we must add the quantity, that is $$\large { h(t) -16(t^2-42t+\square?)+(16\times \square?)\\ \text{so}\\ -(16\times \square?)+(16\times \square?)=0 } $$

OpenStudy (jdoe0001):

a perfect trinomial square has the form of \(\large a^2\pm2ab+b^2\) so our 2nd term, -42t is really => 2(ab), but we already know our "a = t" so -42t => -2(t)(21) so our 3rd number will be \(21^2\)

OpenStudy (jdoe0001):

so $$ h(t) -16(t^2-42t+21^2)+(16\times 21^2)\\ h(t)-16(t-\color{red}{21})^2+\color{red}{7056}\\ \text{so your vertex is at } (\color{red}{21}, \color{red}{7056})\\ $$ what does that mean? well, after 21secs, the projectile is at 7056 feet